OpenStudy (anonymous):

Use fundamental identities to simplify the expression secØ (sinØ/tanØ). I don't remember learning this. please show work, thanks!

OpenStudy (anonymous):

There aren't any inverse functions here?

OpenStudy (anonymous):

oops, nevermind.

OpenStudy (anonymous):

\[\sec(x)=\frac{1}{\sin(x)}\] and\[\tan(x)=\frac{\sin(x)}{\cos(x)}\]replace and then do algebra

OpenStudy (anonymous):

It's a multiple choice, sorry I thought the rest of the question posted. It says, determine which of the following is NOT equivalent. cos^2Ø-sin^2Ø csc^2Ø -cot^2Ø 1 sec^2Ø -tan^2Ø sin^2Ø +cos^2Ø